OpenStudy (anonymous):

find the real solution of x^6+7x^3-8 help

OpenStudy (akshay_budhkar):

x^6+8x^3-x^3-8 X^3(x^3+8)-1(x^3+8) (x^3+8)(x^3-1)

OpenStudy (akshay_budhkar):

if this is 0 then x^3=-8 or x^3=1 therefore x=-2 or x-1

OpenStudy (akshay_budhkar):

get it?

OpenStudy (anonymous):

i don't ghe first pr how did u get +8x^3and all that?

OpenStudy (akshay_budhkar):

see in the first step i split the middle term so as to get suitable factors... you got that step?

OpenStudy (akshay_budhkar):

see the last term -8=(8)(-1) so I wrote 7 as the sum of the factors 7=8-1

OpenStudy (anonymous):

oh i get it :) thank u so much !

OpenStudy (akshay_budhkar):

You are welcome... If you have any more query on it just post here

OpenStudy (akshay_budhkar):

And yes give me my medal lol

OpenStudy (akshay_budhkar):

like this.. you see the option of a good answer just click on it

OpenStudy (akshay_budhkar):

(if you liked it that is hehe)

OpenStudy (anonymous):

hahahah sorry i was writing down the answe hahahah